The EPA says the value of human life is $7.6 million, measured from birth to death. If life expectancy is 78 years, what is the value of the remaining life of an 18-year-old person?
Answer #1

Value of life per year = 7600000 / 78 = 97435.90

The remaining life of the person = 78 - 18 = 60

Total value of remianing life of person = 60 * 97435.90 = 5846153.85